Solve the system: (1/2)x-(1/6)y=10, (2/5)x+(1/2)y=8. Write your answer as aan ordered pair.

Answer by checkley77(12844)

(1/2)x-(1/6)y=10 or x/2-y/6=10
(2/5)x+(1/2)y=8 or 2x/5+y/2=8 multiply this equation by 1/3 & add them.
2x/5*1/3+y/2*1/3=8/3
2x/15+y/6=8/3
x/2-y/6=10
--------------
2x/15+x/2=8/3+10
(2*2x+15x)/30=(8+30)/3
(4x+15x)/30=38/3
19x/30=38/3
3*19x=30*38
57x=1140
x=1140/57
x=20 ans.
20/2-y/6=10
10-y/6=10
10-y/6=10
-y/6=10-10
-y/6=0
-y=6*0
-y=0 ans.
Proof:
2*20/5+y/2=8
40/5+y/2=8
8+y/2=8
y/2=8-8
y/2=0
y=2*0
y=0